ffi aplockuasemaclockmnt_{l}cknexitpollwaitslpinitwirepdainitsplockschgpdaflagsfindenproc_{f}indenproccb_{w}aitdu_{c}onpolldobootdutycdrvsetmustrunactionlistnactionsactioninitsemahisthistposhistmaxsemalistsemainfoinitmonitorpmonitorqmonitor gsc_{1}991.logλnull.logt∂gsc_{1}991.slides gsc_{1}991.slides